# Relational Observables in Group Field Theory

### v1

@@ 1 Introduction
The apparent tension between background independence and locality can be addressed by recognizing that gauge-invariant local observables can be constructed by describing dynamical quantities relative to one another [2]. These relational observables are localized with respect to dynamical, rather than external, reference frames. In the context of classical gravity, such observables were first introduced in [3, 4, 5] and [6, 7, 8, 9], and later developed into more systematic and general constructions in [10, 11, 12, 13, 14, 15, 16, 2], showing in particular their covariance and their equivalence [17, 2, 18] to the “dressed observables” [19, 20, 21, 22] frequently used in gauge theories. The use of relational observables has also been explored in cosmology [23, 24, 25, 26, 27, 28, 29, 30, 31]. At the quantum level, significant progress has been made in the implementation of the relational strategy within (constrained) quantum mechanical systems [32, 33, 34, 35, 36, 37, 38, 39, 40], showing in particular the quantum covariance of physical properties under changes of quantum reference frames (QRFs) [32, 33, 35, 36, 37, 38], as well as the dependence on QRFs of quantum correlations and (thermo)dynamical properties of subsystems [40]. Moreover, recently, an analogous relational description has been implemented within parametrized quantum field theory by using (ideal) QRFs [41]. Despite this progress, a systematic implementation of the relational strategy and study of QRFs within the context of quantum field theory (QFT), even on a fixed background, is still lacking.
@@
In this work, we aim to overcome these limitations by constructing relational observables in GFT that: (i) allow for a clear physical interpretation as observables conditioned on quantum reference frame configurations; (ii) act on the perspective-neutral GFT Fock space; (iii) are defined independently of the quantum state; and (iv) coincide with previously defined relational observables when both are averaged over appropriate collective states, thereby ensuring that they can be successfully used to extract continuum physics, and in particular make contact with previous results in the cosmological setting. To achieve this, we introduce covariant positive operator-valued measures (POVMs) [83, 84, 85] associated with minimally-coupled massless scalar field QRFs (here and throughout the paper, massless is understood to mean without any potential). Following the standard QRF approach [32, 33, 38] (see also [39, 86]), we then use these POVMs to construct relational observables by (on-shell) conditioning perspective-neutral observables on frame orientations, a construction performed here for, to the best of our knowledge, the first time in full QG.
@@ 2 Group Field Theory
where $D^{j}_{mn}(g)$ are Wigner functions, $d(j)=2j+1$, and $\mathcal{I}^{j_{1},\dots,j_{4};\iota}_{n_{1},\dots,n_{4}}$ is an $\mathrm{SU}(2)$ intertwiner that appears due to the right-invariance imposed by equation (1). It is precisely due to the choice of $\mathrm{SU}(2)^{4}$ as the gravitational part of the group domain together with equation (1) that the field modes are decorated with spin network data $\vec{\kappa}\equiv\{\vec{j},\vec{m},\iota\}$, where $\vec{j}$ and $\vec{m}$ are respectively spin and angular momentum labels, while $\iota$ is an intertwiner quantum number.
@@ 2.1 The perspective neutral GFT Fock space
Although historically first introduced in a path-integral formulation (due to the connection with spinfoam models and simplicial gravity explained above) [94], GFT can also naturally be formulated in a canonical language [91, 95]; in particular, the path-integral and the Fock representation of GFT outlined below have been recently shown to be equivalent for a GFT formulation of the Husain-Kuckař theory [96], further confirming that the GFT Fock representation is naturally associated with a free theory—a result consistent with standard QFTs (see [97] for a possible construction in the interacting case). To achieve this reformulation, creation $\hat{\varphi}^{\dagger}(\vec{g},\boldsymbol{\chi})$ and annihilation $\hat{\varphi}(\vec{g},\boldsymbol{\chi})$ operators are introduced that satisfy the following commutation relations:
@@
As in standard QFT, a Fock space $\mathcal{F}$ can be constructed by the repeated action of the creation operator $\hat{\varphi}^{\dagger}_{\vec{\kappa}}(\boldsymbol{\chi})$ on the Fock vacuum |$\ket{0}$. We refer to $\mathcal{F}$ as a “perspective neutral” Fock space, since equations (6) and (7) do not admit a preferred choice of a physical frame, cfr. equation (68). The term “perspective neutral” is used here specifically to highlight that the physical interpretation of $\mathcal{F}$ aligns with that of the perspective neutral space used in the context of QRFs [36, 35, 98, 32, 33, 99, 37, 100, 38], even though the technical construction differs due to the more subtle appearance of diffeomorphism constraints in GFTs (see Sec. 3.3). The perspective-neutral nature of $\mathcal{F}$ is further confirmed by the ability to derive perspective-reduced structures from it, as shown explicitly in Sec. 4.3.
@@
|$\ket{\vec{\kappa};\boldsymbol{\chi}}\equiv\hat{\varphi}^{\dagger}_{\vec{\kappa}}(\boldsymbol{\chi})\ket{0},$ (8)
which represents an open spin-network vertex with spin labels $\vec{\kappa}=(\iota,j_{i},m_{i})$ and scalar field data $\boldsymbol{\chi}=(\chi^{1},\dots,\chi^{m})$. Generic $n$-particle states correspond to a collection of $n$ indistinguishable spin-network vertices. To make the connection with canonical LQG, an important subspace of the Fock space is the space where the vertices are connected via entangled links (specifically, the $m$ labels for each pair of two entangled/connected links form a singlet state [91, 101]); these states can be identified with a graph, and therefore with a spin-network when taking into account the remaining $j$ and $\iota$ $SU(2)$ labels. (Note however that there is a difference between this subspace of the GFT Fock space and the Hilbert space of canonical LQG, as in canonical LQG the spin-networks are embedded in a three-dimensional manifold, in contrast to the abstract graphs that arise in GFT.) As a result, many states |$\ket{\psi}\in\mathcal{F}$ do not correspond to a connected simplicial lattice—it is only states with the entanglement necessary to be fully connected that satisfy this property [91]. Therefore, from the perspective of GFT, the recovery of continuum quantities is a highly non-trivial process, possibly associated with emergent behavior.
@@
|$\ket{\sigma}=\mathcal{N}_{\sigma}\exp\left[\sum_{\vec{\kappa}}\int\mathop{}\!\mathrm{d}\boldsymbol{\chi}\,\sigma_{\vec{\kappa}}(\boldsymbol{\chi})\hat{\varphi}^{\dagger}_{\vec{\kappa}}(\boldsymbol{\chi})\right]|0\rangle\,,\quad|\mathcal{N}_{\sigma}|^{2}=\exp\left[-\sum_{\vec{\kappa}}\int\mathop{}\!\mathrm{d}\boldsymbol{\chi}\,|\sigma_{\vec{\kappa}}(\boldsymbol{\chi})|^{2}\right],$ (11)
these are eigenstates of the annihilation operator $\hat{\varphi}_{\vec{\kappa}(\chi)}$, i.e., $\hat{\varphi}_{\vec{\kappa}(\chi)}\ket{\sigma}=\sigma_{\kappa}(\chi)\ket{\sigma}$.
@@ 3.1 Relational observables
where $f_{*}$ represents the push-forward of $f$. In other words, the frame field provides a map between a kinematical (or coordinate) notion of locality and a physical one [2]. As a simple example, if a scalar field $\chi$ is used as a frame field, $R^{-1}$ is the map from the region $\mathcal{N}\in\mathcal{M}$ where $\chi=\chi_{0}$ to the slice in the configuration space for $\chi$ where $\chi=\chi_{0}$.
@@
×$\quad\qquad\times\left\{\dots\left\{O(\mathbf{x}_{1}),\tilde{C}_{a_{1}}(\mathbf{x}_{1})\right\}\dots,\tilde{C}_{a_{n}}(\mathbf{x}_{n})\right\},$ (15)
[…]
At the quantum level, significant progress has been made for mechanical systems, for any unimodular gauge symmetry group $G$ [38]. The relevant Hilbert space, as described from the perspective of an external frame, is assumed to be of the form $\mathcal{H}_{\text{kin}}=\mathcal{H}_{R}\otimes\mathcal{H}_{S}$, thus factorizing into a quantum reference frame Hilbert space $\mathcal{H}_{R}$ and a system one, $\mathcal{H}_{S}$. Moreover, one assumes that $\mathcal{H}_{\text{kin}}$ carries a unitary tensor product representation of the symmetry group $G$: $\mathcal{H}_{\text{kin}}\ni\ket{\psi}_{\text{kin}}\mapsto U_{RS}(h)\ket{\psi}_{\text{kin}}$.
In this context, the orientations of the quantum reference frame can be represented by $G$-coherent states |$\ket{\phi(g)}_{R}$. Denoting the set of linear operators on the Hilbert space $\mathcal{H}$ by $\mathcal{L}(\mathcal{H})$, then for any operator $\hat{O}^{(S)}\in\mathcal{L}(\mathcal{H}_{S})$, the corresponding relational observable is
$\hat{O}^{(S)}_{R}=\mathcal{G}\left(\ket{\phi(g)}_{R}\bra{\phi(g)}_{R}\otimes\hat{O}^{(S)}\right),$ (16)
@@
is an incoherent group-averaging. In other words, the relational observable $O_{R}^{(S)}$ is constructed by first tensoring the kinematical system observable $O^{(S)}$ with the projector |$\ket{\phi(g)}_{R}\bra{\phi(g)}_{R}$ onto the orientation state of the quantum reference frame (called frame orientation conditioning), and then performing the $G$-twirl that guarantees that the relational observable $O_{S}^{(S)}$ is gauge-invariant.
@@ 3.2 POVMs in QFT and GFT
    $\hat{E}(\cup_{i}X_{i})=\sum_{i}\hat{E}(X_{i})$ for all disjoint sequences $(X_{i})\subset\Sigma$.
These properties ensure (the expectation values of) the POVMs can be interpreted as probabilities; note that $\Omega$ is to be understood as the space containing all possible outcomes of the measurement of interest, performed on any state $|\psi\rangle$ in the Hilbert space $\mathcal{H}$. Operators $\hat{E}(X)$ in the range of POVMs are called effects [83, 39]. If the POVM also satisfies
@@
then the POVM is said to be normalized, and is understood to cover all possible measurement outcomes. Furthermore, let $G$ be a locally compact second countable topological group, .$:G\times\Omega\to\Omega$ a continuous transitive action (so that $\Omega$ is a homogeneous $G$-space) and $U:G\mapsto\mathcal{L}(H)$ a strongly continuous projective unitary representation. Then, a POVM $\hat{E}:\Sigma\to\mathcal{L}(\mathcal{H})$ is called $G$-covariant if it satisfies [83, 32, 38, 39]
@@
As a concrete example of the above definition, the orientation states |$\ket{\phi(g)}_{R}$ used in (16) provide a resolution of the identity on $\mathcal{H}_{R}$, and thus give rise to the following POVM
$\hat{E}_{\phi}(Y)=\int_{Y}\mathop{}\!\mathrm{d}g\,\ket{\phi(g)}\bra{\phi(g)}_{R},$ (18)
@@
Event POVMs in Minkowski spacetime have been investigated in QFT to achieve a consistent quantization of spacetime coordinates [112, 113, 114, 115], particularly time [116, 83, 117, 118]. For this definition to be consistent, event POVMs must transform covariantly under the Poincaré group, see condition (iv) in the definition of POVMs given above. It is also important to note that, unlike standard POVMs, event POVMs in QFT cannot be normalized, as the vacuum state cannot define any event [113].
@@
First, we introduce the effect density
$\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ $=\mathop{}\!\mathrm{d}\chi\sum_{n=1}^{\infty}\frac{1}{n!}\sum_{\vec{\kappa}_{1},\dots\vec{\kappa}_{n}}\!\int\!\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]\frac{\sum_{i=1}^{n}\delta(\chi_{i}-\chi)}{n}\left[\prod_{i=1}^{n}\hat{\varphi}_{\vec{\kappa}_{i}}^{\dagger}(\chi_{i})\right]\!\ket{0}\bra{0}\!\left[\prod_{i=1}^{n}\hat{\varphi}_{\vec{\kappa}_{i}}(\chi_{i})\right]$
@@
$\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\equiv\frac{1}{n!}\left[\prod_{i=1}^{n}\hat{\varphi}^{\dagger}_{\vec{\kappa}_{i}}(\chi_{i})\right]\ket{0}\bra{0}\left[\prod_{i=1}^{n}\hat{\varphi}_{\vec{\kappa}_{i}}(\chi_{i})\right],$ (21)
@@
×$\quad\times\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\,,$ (22)
@@
where $X$ is any subspace of $\Omega=\mathbb{R}$, namely (possibly disjoint unions of) intervals of $\mathbb{R}$. An important simplification arises due to the fact that the properties of the density $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ are satisfied by construction also by the operator $\hat{E}_{\chi}(X)$. For this reason, from here on, we will focus exclusively on the density (3.2).
@@
|$\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}=\frac{1}{n!}\prod_{i=1}^{n}\hat{\varphi}_{\vec{\kappa}_{i}}^{\dagger}(\chi_{i})\ket{0}.$ (25)
@@
$\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}=\mathop{}\!\mathrm{d}\chi\sum_{i=1}^{n}\frac{\delta(\chi_{i}-\chi)}{n}\ket{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}},$ (26)
Thus, for each $n$-body sector of the GFT Fock space (except for the vacuum which it annihilates, $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\ket{0}=0$), $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ localizes one of the GFT quanta at the clock value $\chi$. From equation (26), together with the fact that the Fock vacuum combined with the $n$-particle states provide a basis for the GFT Fock space, we see that $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\geq 0$. This, together with the $\sigma$-additivity of the measure $\mathop{}\!\mathrm{d}\chi$ implies that the effect operators (23) do indeed constitute a POVM.
@@
|$\ket{0}\bra{0}+\int_{\mathbb{R}}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)=\mathbbm{1}_{\mathcal{F}}\,.$ (27)
@@
Crucially, in the same way as the first moment of event POVMs define coordinate observables, the first moment of the density (3.2) gives the observable

### v2

@@ 1 Introduction
The apparent tension between background independence and locality can be addressed by recognizing that gauge-invariant local observables can be constructed by describing dynamical quantities relative to one another [2]. These relational observables are localized with respect to dynamical, rather than external, reference frames. In the context of classical gravity, such observables were first introduced in [3, 4, 5] and [6, 7, 8, 9], and later developed into more systematic and general constructions in [10, 11, 12, 13, 14, 15, 16, 2], showing in particular their covariance and their equivalence [17, 2, 18] to the “dressed observables” [19, 20, 21, 22] frequently used in gauge theories. The use of relational observables has also been explored in cosmology [23, 24, 25, 26, 27, 28, 29, 30, 31]. At the quantum level, significant progress has been made in the implementation of the relational strategy within (constrained) quantum mechanical systems [32, 33, 41, 34, 35, 36, 37, 38, 39, 40, 42], showing in particular the quantum covariance of physical properties under changes of quantum reference frames (QRFs) [32, 33, 34, 35, 36, 37, 43, 44], as well as the dependence on QRFs of quantum correlations and (thermo)dynamical properties of subsystems [39, 45]. Moreover, recently, an analogous relational description has been implemented within parametrized quantum field theory by using (ideal) QRFs [46]. Despite this progress, a systematic implementation of the relational strategy and study of QRFs within the context of quantum field theory (QFT), even on a fixed background, is still lacking.
@@
In this work, we aim to overcome these limitations by adapting to the GFT formalism the QRF approach of building relational observables by conditioning system observables on the orientation of a reference frame. In this way, we can construct relational observables in GFT that: (i) allow for a clear physical interpretation as observables conditioned on quantum reference frame configurations; (ii) act on the perspective-neutral GFT Fock space; (iii) are defined independently of the quantum state; and (iv) coincide with previously defined relational observables when both are averaged over appropriate collective states, thereby ensuring that they can be successfully used to extract continuum physics, and in particular make contact with previous results in the cosmological setting. More specifically, to achieve this we introduce covariant positive operator-valued measures (POVMs) [88, 89, 90] associated with minimally-coupled massless scalar field QRFs (here and throughout the paper, massless is understood to mean without any potential). Motivated by the standard QRF approach [32, 33, 41, 34, 35, 36, 37, 38, 39, 40, 42] (see also [91]), we then use these POVMs to construct relational observables by (on-shell) conditioning perspective-neutral observables on frame orientations, a construction performed here for, to the best of our knowledge, the first time in full QG.
@@ 2 Group Field Theory
where $D^{j}_{mn}(g)$ are Wigner rotation matrices, $d(j)=2j+1$, and $\mathcal{I}^{j_{1},\dots,j_{4};\iota}_{n_{1},\dots,n_{4}}$ is an $\mathrm{SU}(2)$ intertwiner that appears due to the right-invariance imposed by equation (1). It is precisely due to the choice of $\mathrm{SU}(2)^{4}$ as the gravitational part of the group domain together with equation (1) that the field modes are decorated with spin network data $\vec{\kappa}\equiv\{\vec{j},\vec{m},\iota\}$, where $\vec{j}$ and $\vec{m}$ are respectively spin and angular momentum labels, while $\iota$ is an intertwiner quantum number.
@@ 2.1 The perspective neutral GFT Fock space
Although historically first introduced in a path-integral formulation (due to the connection with spinfoam models and simplicial gravity explained above) [100], GFT can also naturally be formulated in a canonical language [96, 101]; in particular, the path-integral and the Fock representation of GFT outlined below have been recently shown to be equivalent for a GFT formulation of the Husain-Kuchař theory [102], further confirming that the GFT Fock representation is naturally associated with a free theory—a result consistent with standard QFTs (the proof of this equivalence is possible since the GFT is non-interacting, see [103] for a possible construction in the interacting case). To achieve this reformulation, creation $\hat{\varphi}^{\dagger}(\vec{g},\boldsymbol{\chi})$ and annihilation $\hat{\varphi}(\vec{g},\boldsymbol{\chi})$ operators are introduced that satisfy the following commutation relations:
@@
As in standard QFT, a Fock space $\mathcal{F}$ can be constructed by the repeated action of the creation operator $\hat{\varphi}^{\dagger}_{\vec{\kappa}}(\boldsymbol{\chi})$ on the Fock vacuum $|{0}\rangle$. We refer to $\mathcal{F}$ as a “perspective neutral” Fock space, since equations (6) and (7) do not admit a preferred choice of a physical frame, cfr. equation (68). The term “perspective neutral” is used here to highlight that the physical interpretation of $\mathcal{F}$ is analogous to that of the perspective neutral space used in the context of QRFs [35, 34, 104, 32, 33, 105, 36, 106, 37], even though the technical construction differs for two important reasons.
A first reason regards the imposition of gauge-invariance. In the language of QRFs, the relevant symmetry group for quantum gravity would be (small) diffeomorphisms. In the GFT quantization of gravity, the continuum manifold associated with this symmetry is eliminated entirely, and the theory is instead defined on a domain given by the possible values (images) of the fields (both matter and gravitational). As a result, diffeomorphism invariance cannot be imposed directly as a symmetry of the theory; rather, the consequences of this classical symmetry are encoded in the intrinsically background-independent nature of the GFT quantization. In this sense, the GFT Fock space is not obtained through the explicit imposition of a gauge constraint, as is typically the case in QRF frameworks. However, analogously to the QRF case, this space encodes and links all possible internal frame perspectives. Indeed, all fields that may be used as internal frames are included in the GFT field domain and are treated on an equal footing in equations (6) and (7). By contrast, equation (68) involves “same-time” commutation relations that induce a Fock space structure adapted to the perspective of a specific frame—more precisely, a chosen clock. This is the reason why we refer to the former as “perspective-neutral” and the latter as “perspective-reduced.” The perspective-neutral nature of $\mathcal{F}$ is further confirmed by the ability to derive perspective-reduced structures from it, as shown explicitly in Sec. 4.3.
A second point concerns the Hilbert space structure of the GFT Fock space. While a distinction between “frame” and “system” degrees of freedom can be made at the level of the GFT field domain—where classical fields can be clearly separated due to the factorized structure of the domain—this factorization does not persist at the quantum level. In the full quantum theory, both the frame and system degrees of freedom are expected to emerge collectively from the behavior of the GFT field. Note that this lack of factorization is, in fact, to be expected after the imposition of the quantum constraint. Both these points will be relevant in the construction of relational observables in Sec. 3.3.
@@
$|{\vec{\kappa};\boldsymbol{\chi}}\rangle\equiv\hat{\varphi}^{\dagger}_{\vec{\kappa}}(\boldsymbol{\chi})|{0}\rangle,$ (8)
which represents an open spin-network vertex with spin labels $\vec{\kappa}=(\iota,j_{i},m_{i})$ and scalar field data $\boldsymbol{\chi}=(\chi^{1},\dots,\chi^{m})$. Generic $n$-particle states correspond to a collection of $n$ indistinguishable spin-network vertices. To make the connection with canonical LQG, an important subspace of the Fock space is the space where the vertices are connected via entangled links (specifically, the $m$ labels for each pair of two entangled/connected links form a singlet state [96, 107]); these states can be identified with a graph, and therefore with a spin-network when taking into account the remaining $j$ and $\iota$ $SU(2)$ labels. (Note however that there is a difference between this subspace of the GFT Fock space and the Hilbert space of canonical LQG, as in canonical LQG the spin-networks are embedded in a three-dimensional manifold, in contrast to the abstract graphs that arise in GFT.) As a result, many states $|{\psi}\rangle\in\mathcal{F}$ do not correspond to a connected simplicial lattice—it is only states with the entanglement necessary to be fully connected that satisfy this property [96] (see also [97] for a concrete construction of some GFT states that can be associated with corresponding LQG kinematical states). Therefore, from the perspective of GFT, the recovery of continuum quantities is a highly non-trivial process, possibly associated with emergent behavior.
@@
$|{\sigma}\rangle=\mathcal{N}_{\sigma}\exp\left[\sum_{\vec{\kappa}}\int\mathop{}\!\mathrm{d}\boldsymbol{\chi}\,\sigma_{\vec{\kappa}}(\boldsymbol{\chi})\hat{\varphi}^{\dagger}_{\vec{\kappa}}(\boldsymbol{\chi})\right]|0\rangle\,,\quad|\mathcal{N}_{\sigma}|^{2}=\exp\left[-\sum_{\vec{\kappa}}\int\mathop{}\!\mathrm{d}\boldsymbol{\chi}\,|\sigma_{\vec{\kappa}}(\boldsymbol{\chi})|^{2}\right],$ (11)
these are eigenstates of the annihilation operator $\hat{\varphi}_{\vec{\kappa}(\chi)}$, i.e., $\hat{\varphi}_{\vec{\kappa}(\chi)}|{\sigma}\rangle=\sigma_{\kappa}(\chi)|{\sigma}\rangle$.
@@ 3.1 Relational observables
where $f_{*}$ represents the push-forward of $f$. In other words, the frame field provides a map between a kinematical (or coordinate) notion of locality and a physical one [2]. As a simple example, if a scalar field $\chi$ is used as a frame field, $R^{-1}$ maps any slice $\chi=\chi_{0}$ in the manifold to the corresponding subset in configuration space.
@@
$\quad\qquad\times\left\{\dots\left\{O(\mathbf{x}_{1}),\tilde{C}_{a_{1}}(\mathbf{x}_{1})\right\}\dots,\tilde{C}_{a_{n}}(\mathbf{x}_{n})\right\},$ (15)
where $G^{\mu}(\mathbf{x})=\chi^{a}_{0}-\chi^{a}$ are gauge-fixing conditions depending on the frame field $R^{-1}=\chi$, $\tilde{C}_{a}$ are (weakly abelianized) constraints generating the gauge symmetries of the theory, $\mathbf{x}$ are coordinates on the chosen Cauchy slice $\Sigma$, and $\{\cdot,\cdot\}$ are kinematical Poisson brackets [10, 11].
At the quantum level, significant progress has been made for mechanical systems, for any unimodular gauge symmetry group $G$ [37]. The relevant Hilbert space, as described from the perspective of an external frame, is assumed to be of the form $\mathcal{H}_{\text{kin}}=\mathcal{H}_{R}\otimes\mathcal{H}_{S}$, thus factorizing into a quantum reference frame Hilbert space $\mathcal{H}_{R}$ and a system one, $\mathcal{H}_{S}$. Moreover, one assumes that $\mathcal{H}_{\text{kin}}$ carries a unitary tensor product representation of the symmetry group $G$: $\mathcal{H}_{\text{kin}}\ni|{\psi}\rangle_{\text{kin}}\mapsto U_{RS}(h)|{\psi}\rangle_{\text{kin}}=U_{R}(g)\otimes U_{S}(g)|{\psi}\rangle_{\text{kin}}$.
In this context, the orientations of the quantum reference frame can be represented by $G$-coherent states $|{\phi(g)}\rangle_{R}$, generated by the action of the group on a “seed” state $|{\phi(e)}\rangle\in\mathcal{H}_{R}:|{\phi(g)}\rangle=U_{R}(g)|{\phi(e)}\rangle$ [108]. Assuming that $G$ acts transitively and freely on the coherent state system, there is an invertible equivariant map $e:G\to\mathcal{H}_{R}$, so that the above frame orientations can completely parametrize the gauge orbits in the Hilbert space or observable algebra [37]. Denoting the set of linear operators on the Hilbert space $\mathcal{H}$ by $\mathcal{L}(\mathcal{H})$, then for any operator $\hat{O}^{(S)}\in\mathcal{L}(\mathcal{H}_{S})$, the corresponding relational observable is [37] (and see also [40, 41, 38, 109])
$\hat{O}^{(S)}_{R}=\mathcal{G}\left(|{\phi(g)}\rangle_{R}\langle{\phi(g)}|_{R}\otimes\hat{O}^{(S)}\right),$ (16)
@@
is an incoherent group-averaging. In other words, the relational observable $O_{R}^{(S)}$ is constructed by first tensoring the kinematical system observable $O^{(S)}$ with the projector $|{\phi(g)}\rangle_{R}\langle{\phi(g)}|_{R}$ onto the orientation state of the quantum reference frame (called frame orientation conditioning), and then performing the $G$-twirl that guarantees that the relational observable $O_{S}^{(S)}$ is gauge-invariant. Note that the integral in (17) generally does not converge for non-compact groups; see [37] for a more detailed discussion. However, as we will see in Sec. 3.3, the construction of relational observables in GFT does not require such group averaging, thereby avoiding such convergence issues.
@@ 3.2 POVMs in QFT and GFT
    For any countable collection of disjoint sets $X_{i}\subset\Sigma$, $\hat{E}(\cup_{i}X_{i})=\sum_{i}\hat{E}(X_{i})$, (with convergence in the weak operator topology required in the case of an infinite series).
Operators $\hat{E}(X)$ in the range of POVMs are called effects [88, 38]. If the POVM also satisfies
@@
then the POVM is said to be normalized, and is understood to cover all possible measurement outcomes. Together, these three properties ensure (the expectation values of) the POVMs can be interpreted as probabilities; note that $\Omega$ is to be understood as the space containing all possible outcomes of the measurement of interest, performed on any state $|\psi\rangle$ in the Hilbert space $\mathcal{H}$.
Furthermore, let $G$ be a locally compact second countable topological group, $G\times\Omega\to\Omega$ a continuous transitive action (so that $\Omega$ is a homogeneous $G$-space) and $U:G\mapsto\mathcal{L}(H)$ a strongly continuous projective unitary representation. Then, a POVM $\hat{E}:\Sigma\to\mathcal{L}(\mathcal{H})$ is called $G$-covariant if it satisfies [110, 111, 113, 115, 114, 88, 112] (see also [32, 37, 38] for applications in the context of QRFs):
@@
As a concrete example of the above definition, the orientation states $|{\phi(g)}\rangle_{R}$ used in (16) provide a resolution of the identity on $\mathcal{H}_{R}$, and thus give rise to the following POVM
$\hat{E}_{\phi}(Y)=\int_{Y}\mathop{}\!\mathrm{d}g\,|{\phi(g)}\rangle\langle{\phi(g)}|_{R},$ (18)
@@
Event POVMs in Minkowski spacetime have been investigated in QFT to achieve a consistent quantization of spacetime coordinates [126, 127, 128, 129, 130], (time quantization has also been investigated in quantum mechanics, see, e.g., [131, 88, 132, 133]). For this definition to be consistent, event POVMs must transform covariantly under the Poincaré group, see condition (iv) in the definition of POVMs given above. It is also important to note that, unlike standard POVMs, event POVMs in QFT cannot be normalized, as the vacuum state cannot define any event [127].
@@
Different definitions for the effect density are possible, so long as they satisfy the first two properties (i) and (ii) for POVMs; however, a natural choice that has nice properties is the following effect density (for another possible choice, see Sec. 4.3):
$\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ $=\mathop{}\!\mathrm{d}\chi\sum_{n=1}^{\infty}\frac{1}{n!}\sum_{\vec{\kappa}_{1},\dots\vec{\kappa}_{n}}\!\int\!\left[\prod_{i=1}^{n}\mathop{}\!\mathrm{d}\chi_{i}\right]\frac{\sum_{i=1}^{n}\delta(\chi_{i}-\chi)}{n}\left[\prod_{i=1}^{n}\hat{\varphi}_{\vec{\kappa}_{i}}^{\dagger}(\chi_{i})\right]\!|{0}\rangle\langle{0}|\!\left[\prod_{i=1}^{n}\hat{\varphi}_{\vec{\kappa}_{i}}(\chi_{i})\right]$
@@
$\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\equiv\frac{1}{n!}\left[\prod_{i=1}^{n}\hat{\varphi}^{\dagger}_{\vec{\kappa}_{i}}(\chi_{i})\right]|{0}\rangle\langle{0}|\left[\prod_{i=1}^{n}\hat{\varphi}_{\vec{\kappa}_{i}}(\chi_{i})\right],$ (21)
@@
$\quad\times\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\,.$ (22)
@@
where $X$ is any (or possibly disjoint unions of) interval(s) of $\Omega=\mathbb{R}$. An important simplification arises due to the fact that the properties of the density $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ are satisfied by construction also by the operator $\hat{E}_{\chi}(X)$. For this reason, from here on, we will focus exclusively on the density (3.2).
@@
$|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle=\frac{1}{n!}\prod_{i=1}^{n}\hat{\varphi}_{\vec{\kappa}_{i}}^{\dagger}(\chi_{i})|{0}\rangle.$ (25)
@@
$\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle=\mathop{}\!\mathrm{d}\chi\sum_{i=1}^{n}\frac{\delta(\chi_{i}-\chi)}{n}|{\chi_{1},\vec{\kappa}_{1};\dots,\chi_{n},\vec{\kappa}_{n}}\rangle,$ (26)
Since the action of the operator $\hat{E}_{\chi}(d\chi)$ on a generic $n$-particle state is non-zero only if at least one of the particles contained in the state has a clock quantum number equal to $\chi$, for each $n$-body sector of the GFT Fock space (except for the vacuum which it annihilates, $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)|{0}\rangle=0$), $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)$ localizes one of the GFT quanta at the clock value $\chi$. From equation (26), together with the fact that the Fock vacuum combined with the $n$-particle states provide a basis for the GFT Fock space, we see that $\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)\geq 0$. This, together with the $\sigma$-additivity of the measure $\mathop{}\!\mathrm{d}\chi$ implies that the effect operators (23) do indeed constitute a POVM.
@@
$|{0}\rangle\langle{0}|+\int_{\mathbb{R}}\hat{E}_{\chi}(\mathop{}\!\mathrm{d}\chi)=\mathbbm{1}_{\mathcal{F}}\,.$ (27)
@@
Crucially, in the same way as the first moment of event POVMs define coordinate observables, the first moment of the density (3.2) gives the observable (see A.1 for a proof that this operator is essentially self-adjoint)